# Provably Near-Optimal Federated Ensemble Distillation with Negligible Overhead

### v1

@@ 1 Introduction
To address the data heterogeneity issue, various approaches have been taken, including regularizing the objectives of the client models [29, 37, 41, 66, 46], normalizing features or weights [13, 32], utilizing past round models [66, 58], sharing feature information [11, 65, 55], introducing personalized layers [25], and learning the average input-output relation of client models through ensemble distillation [5, 17, 12, 50, 42, 7, 62, 47, 57, 54, 71, 69]. In particular, the last approach, federated ensemble distillation, has recently gained significant attention for its effectiveness in mitigating data heterogeneity and for its advantage of being effectively applicable to heterogeneous client models. It requires an unlabeled dataset at the server, for which pseudo labels are created based on client predictions. By training on this pseudo-labeled dataset at the server, the server distills the knowledge from the clients. This additional dataset can be either public [5, 17, 12, 50], held only by the server due to its exceptional data collection capability [42, 7, 62, 47], or obtained through a data-free approach [57, 54, 71, 69]. Note that the performance of ensemble distillation depends on the quality of the pseudo-labels, which ultimately translates into a problem of appropriately assigning weights to client predictions for each data point, particularly in situations of data heterogeneity. In this research stream of federated ensemble distillation, early studies like FedDF [42] applied uniform weighting. Subsequently, algorithms such as Fed-ET [7], FedHKT [12], FedDS [47], and DaFKD [57] emerged, which utilize metrics like variance, entropy, and judgement of client discriminator as indicators of confidence in client predictions for weighting. However, analysis regarding the rationale behind optimal weighting remains scarce.
@@ 2 System Model and Related Work
FedDF [42], Fed-ET [7] $\mathcal{L}_{p}(\sum_{k=1}^{K}\alpha_{k}h_{\hat{p}_{k}}^{*})\leq\sum_{k=1}^{K}\alpha_{k}\cdot\left[\mathcal{L}_{{\hat{p}_{k}}}(h_{\hat{p}_{k}}^{*})+\frac{1}{2}d_{\mathcal{H}\triangle\mathcal{H}}(p_{k},p)+\lambda_{k}+O\left(\frac{\log(\delta^{-1})}{\sqrt{n_{k}}}\right)\right]$ (1.1)
@@
These bounds relate the loss of an ensemble model (the LHS of (1), (1) and (1)) to the average empirical loss of client models (the first term in the RHS of (1), (1) and (1)). Note that the bound (1) assumes a fixed weight per client irrelevant to data points, hence there is a lack of analysis for assigning varying weights per data point. Furthermore, (1) provides a generalization bound that becomes loose as it scales with the average distribution discrepancy between $p$ and $p_{k}$ across clients. The bound (1) assumes a specific weighting function of each data point, but it is too loose because it becomes vacuous as $K$ increases. In contrast, (1), which is derived in this paper, provides a refined generalization bound for weighting functions. To the best of our knowledge, this is the tightest bound currently provided, as detailed in Theorem C.1.
@@ 4.1 Experimental Setting
For architecture, we employed ResNet-18 [22] with batch normalization layers [27]. For baselines, we considered the vanilla FedAVG [45], FedProx [37], SCAFFOLD [29] and FedDisco [67] that do not perform ensemble distillation, FedDF [42], FedGKD+ [66] and DaFKD [57] that incorporate ensemble distillation. For comparison with other weighting methods, we considered the variance-based weighting method of Cho et al. [7], the entropy-based methods of Deng et al. [12] and Park et al. [47], and the domain-aware method of Wang et al. [57], described in Section 2. As an upper bound of the performance, we also compared with central training that trains the server model directly using the entire client dataset. FedGO and DaFKD require image generators and discriminators. For the generator, we considered the three approaches (G1), (G2), and (G3) in Section 3.2. For (G1) and (G3), we adopted the model architecture and training method proposed in WGAN-GP [20]. For (G2), we employed StyleGAN-XL [51], pretrained on ImageNet [35]. Unless specified otherwise, we assume (G1). For discriminators, we utilized a 4-layer CNN. More experimental details are provided in Appendix E.2.
@@ 4.2 Results
In Appendix F, we provide more experimental results. We report ensemble test accuracy of the baselines and FedGO, demonstrating a larger improvement compared to test accuracy. We also provide results for scenarios involving larger clients or different model structures, cases where the server dataset is different from the client datasets, and data-free approaches when no server dataset is available, showing significant performance gains over the baselines. Additionally, we report the performance of FedGO with a reduced server dataset and various discriminator training epochs, showing that even with only 20% of the server dataset, FedGO achieves a performance gain of 15%p over FedAVG. Furthermore, FedGO outperforms the baselines even with significantly fewer discriminator training epochs.
@@ 5 Conclusion
We proposed the FedGO algorithm, which effectively addresses the challenge of client data heterogeneity. Our algorithm was proposed based on theoretical analysis of optimal ensemble distillation, and various experimental results demonstrated its high performance and fast convergence rate under various scenarios with and without extra server dataset. Due to page limit, limitation and broader impact of our work are provided in Appendices H and I, respectively.

### v2

@@ 1 Introduction
To address the data heterogeneity issue, various approaches have been taken, including regularizing the objectives of the client models [29, 37, 41, 67, 46, 54], normalizing features or weights [13, 32], utilizing past round models [67, 59], sharing feature information [11, 66, 56, 72], introducing personalized layers [25], and learning the average input-output relation of client models through ensemble distillation [5, 17, 12, 50, 42, 7, 63, 47, 58, 55, 73, 70]. In particular, the last approach, federated ensemble distillation, has recently gained significant attention for its effectiveness in mitigating data heterogeneity and for its advantage of being effectively applicable to heterogeneous client models. It requires an unlabeled dataset at the server, for which pseudo labels are created based on client predictions. By training on this pseudo-labeled dataset at the server, the server distills the knowledge from the clients. This additional dataset can be either public [5, 17, 12, 50], held only by the server due to its exceptional data collection capability [42, 7, 63, 47], or obtained through a data-free approach [58, 55, 73, 70]. Note that the performance of ensemble distillation depends on the quality of the pseudo-labels, which ultimately translates into a problem of appropriately assigning weights to client predictions for each data point, particularly in situations of data heterogeneity. In this research stream of federated ensemble distillation, early studies like FedDF [42] applied uniform weighting. Subsequently, algorithms such as Fed-ET [7], FedHKT [12], FedDS [47], and DaFKD [58] emerged, which utilize metrics like variance, entropy, and judgement of client discriminator as indicators of confidence in client predictions for weighting. However, analysis regarding the rationale behind optimal weighting remains scarce.
@@
For ease of reproduction, our code is open-sourced (https://github.com/pupiu45/FedGO).
@@ 2 System Model and Related Work
[…]
@@
These bounds relate the loss of an ensemble model (the LHS of (1), (1) and (1)) to the average empirical loss of client models (the first term in the RHS of (1), (1) and (1)). Note that the bound (1) assumes a fixed weight per client irrelevant to data points, hence there is a lack of analysis for assigning varying weights per data point. Furthermore, (1) provides a generalization bound that becomes loose as it scales with the average distribution discrepancy between $p$ and $p_{k}$ across clients. The bound (1) assumes a specific weighting function of each data point, but it is too loose because it becomes vacuous as $K$ increases. In contrast, (1), which is derived in this paper, provides a refined generalization bound for weighting functions. To the best of our knowledge, this is the tightest bound currently provided, as detailed in Theorem C.1. Furthermore, we also emphasize that the bound (1) is independent of client data heterogeneity, as it does not include any client data discrepancy term.
@@ 4.1 Experimental Setting
For architecture, we employed ResNet-18 [22] with batch normalization layers [27]. For baselines, we considered the vanilla FedAVG [45], FedProx [37], SCAFFOLD [29], FedDisco [68], FedUV [54] and FedTGP [72] that do not perform ensemble distillation, FedDF [42], FedGKD+ [67] and DaFKD [58] that incorporate ensemble distillation. For comparison with other weighting methods, we considered the variance-based weighting method of Cho et al. [7], the entropy-based methods of Deng et al. [12] and Park et al. [47], and the domain-aware method of Wang et al. [58], described in Section 2. As an upper bound of the performance, we also compared with central training that trains the server model directly using the entire client dataset. FedGO and DaFKD require image generators and discriminators. For the generator, we considered the three approaches (G1), (G2), and (G3) in Section 3.2. For (G1) and (G3), we adopted the model architecture and training method proposed in WGAN-GP [20]. For (G2), we employed StyleGAN-XL [51], pretrained on ImageNet [35]. Unless specified otherwise, we assume (G1). For discriminators, we utilized a 4-layer CNN. More experimental details are provided in Appendix E.2.
@@ 4.2 Results
In Appendix F, we provide more experimental results. We report ensemble test accuracy of the baselines and FedGO, demonstrating a larger improvement compared to test accuracy. We also provide results for scenarios involving larger clients or different model structures, cases where the server dataset is different from the client datasets, and data-free approaches when no server dataset is available, showing significant performance gains over the baselines. Additionally, we report the performance of FedGO with a reduced server dataset and various discriminator training epochs, showing that even with only 20% of the server dataset, FedGO achieves a performance gain of 15%p over FedAVG. Furthermore, FedGO outperforms the baselines even with significantly fewer discriminator training epochs. Finally, we demonstrate the robustness of FedGO under adversarial conditions where a subset of clients behave in a Byzantine manner, showing that our method maintains superior performance even when up to 50% of the client discriminators are compromised.
@@ 5 Conclusion
We proposed the FedGO algorithm, which effectively addresses the challenge of client data heterogeneity. Our algorithm was proposed based on theoretical analysis of optimal ensemble distillation, and various experimental results demonstrated its high performance and fast convergence rate under various scenarios with and without extra server dataset. The limitation of our work is provided in Appendix H.
Acknowledgements
This work was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) through the Next Generation Semantic Communication Network Research Center Grant (RS-2024-00398948) funded by the Korean Government (MSIT), and in part by the IITP through 6G $\cdot$ Cloud Research and Education Open Hub Grant (RS-2024-00428780) funded by the Korea government (MSIT).
Impact Statement
@@ E.2 Detailed Experimental Settings for Image Classification Tasks
FedUV [54] introduces two regularization terms—classifier variance and encoder representation uniformity—to emulate IID conditions in federated learning, mitigating local bias in non-IID settings. The loss function combines the standard classification loss with the uniformity term (weighted by $\lambda$) and the variance term (weighted by $\alpha$). By following Son et al. [54], we set the hyperparameters $\lambda=0.5$ and $\alpha=2.5$ as they achieved the best trade-off between generalization and personalization across clients.
FedTGP [72] proposes Trainable Global Prototypes (TGP) with Adaptive-Margin-Enhanced Contrastive Learning to address data heterogeneity by improving class separability and semantic alignment among clients. The model trains prototypes using contrastive learning with a temperature-scaled margin. By following Zhang et al. [72], for the TGP model, we used a batch size of 10, trained for 1,000 epochs with a learning rate of 0.005 and a margin parameter $\tau$ = 100. The contrastive loss was weighted by $\lambda$ = 0.001 to balance its influence with the main task objective.
@@
FedTGP [72] utilizes an additional TGP model. We implemented TGP model following its official open-source implementation from https://github.com/TsingZ0/FedTGP.
@@ F.7.3 Discriminator Architectures
F.8 Impact of Byzantine Clients
We have conducted additional experiments in which 5 and 10 out of 20 clients were Byzantine, outputting only the maximum value for the discriminator. As shown in Table 15, the classification accuracy on CIFAR-10 without any Byzantine clients was 72.35% with a standard deviation of 9.01. When 5 clients were Byzantine, the accuracy dropped to 69.75%, and further decreased to 66.38% when 10 clients were Byzantine. Despite this challenging scenario where up to half of the clients were compromised, our method still significantly outperformed all baseline approaches that do not incorporate a discriminator, demonstrating its strong robustness against adversarial behavior.
Table 15: Server test accuracy (%) of FedGO on CIFAR-10 with $\alpha=0.05$ at the 100-th communication round along with the number of Byzantine clients.
Byzantine Clients Accuracy (%)
0 72.35$\pm$9.01
5 69.75$\pm$5.05
10 66.38$\pm$4.97